Complete the following proof:
1.
B is between A and D. C is between B and D.
2.
AD = AB + BD
3.
BD = BC + CD
4.
AD = AB + BC + CD

Answer by MathLover1(20850)


1.
B is between A and D. C is between B and D.
2.
AD = AB + BD
3.
BD = BC + CD
4.
AD = AB + BC + CD
5.
AC +CD=AD....................segment addition property
6.
AD =AC +CD....................segment property
7.
AB + BD=AC +CD......................transitive property